OpenStudy (anonymous):

The price of a large pizza increased from $10 to $12. What was the percent of increase in the cost of the pizza? Round to the nearest percent, if necessary. A. 20% B. 83% C. 17% D. 2%

OpenStudy (triciaal):

what is the change in price new - old = 12 - 10 = 2 how do you express percent per 100 (2/10)*100 = 20% increase